# Dynamic instabilities and turbulence of merged rotating Bose-Einstein condensates

### v1

@@ III.2 Effect of rotation on the dynamics and turbulence state of the merged condensate
Figure 13: Evolution of different components of condensate energies with respect to time for (a) $t_{m}=4$, (b) $t_{m}=6$, (c) $t_{m}=10$, and (d) $t_{m}=20$ for rotating case with $\Omega=0.45$.
@@
Similar to the non-rotating case, all energy components (except for rotation energy) exhibit oscillations consistent with the trap frequency for shorter merging times. However, unlike the non-rotating case, the amplitude of these oscillations damps out quickly [cf. Figs. 13(b)-(d)]. The kinetic energy decreases with increasing merging time, but unlike the non-rotating merger, it does not go to zero for the $t_{m}=10$ and $t_{m}=20$ cases.
@@
Figure 14: Evolution of chemical potential (averaged over $3000$ time units from the time of merger) of the merged condensate with respect to different merging times for (a) $\Omega=0.0$ (black line with empty squares) and $\Omega=0.2$, (red line with empty circles) (b) $\Omega=0.3$ (green line with empty triangles) and $\Omega=0.45$ (blue line with empty triangles) frequencies, respectively. The dashed horizontal lines represent the ground state chemical potential for the merged condensate rotating at $2\Omega$.
@@
Figure 15: Time-averaged incompressible kinetic energy spectra for the rotating case ($\Omega=0.45$), for merging time $t_{m}=4$ in the time range (a) $t=4$ to $t=24$ (b) $t=24$ to $t=44$ and for merging time $t_{m}=6$ in the time range (c) $t=6$ to $t=26$ and (d) $t=26$ to $t=46$. Both cases initially showing $k^{-5/3}$ and later a $k^{-1}$ scaling.
@@
Achieving an adiabatic merger in the condensate depends on both the merging time ($t_{m}$) and the rotation frequency ($\Omega$), as shown in Fig. 14. Condensates rotating at frequency $\Omega$ merge to form a condensate rotating at $2\Omega$. Sufficiently long merging times ensure that this merged condensate reaches its ground state. Notably, for both non-rotating condensates and those at low rotation frequencies [black line with empty squares in Fig. 14(a)], the ground state is achieved with merging times of $t_{m}\sim 20$. In Figs. 14(b), as rotation frequency increases, longer merging times are required for the final condensate to approach its ground state. Even without soliton decay, trap movement can create compressible excitations. These excitations persist longer at higher rotation frequencies, as rotational forcing prevents them from dissipating, thus requiring much slower mergers to suppress them.
@@
The incompressible kinetic energy spectra for the rotating case exhibit similar scaling behaviour to that of the non-rotating case (see Fig. 15). It is important to note that a forcing peak emerges due to rotation at $k<2\pi/R$ gradually diminishes for longer merging times. For cases with $t_{m}=4$ and $t_{m}=6$, a short-lived $k^{-5/3}$ scaling is initially observed, which later transforms into a $k^{-1}$ scaling, corresponding to the formation of vortex-rings (see Ref. [58]). These vortex rings are formed in the soliton decay process through snake instability, as illustrated in Ref. [51].
@@
In contrast to the non-rotating case, where no scaling is observed for $t_{m}=10$ and $t_{m}=20$, the rotating case exhibits $k^{-5/3}$ scaling, albeit for a relatively shorter $k$ range initially. However, the $k^{-1}$ scaling persists for the $t_{m}=10$ case but not for the $t_{m}=20$ case, as no soliton is formed to decay into vortex-rings. In the non-rotating scenario, the scaling behaviour of the incompressible spectra disappears due to vortex annihilation. Also, in contrast to the non-rotating case, the $k^{-5/3}$ scaling appears consistently for length scales larger than inter-vortex distance $\ell_{0}$, where collective vortex dynamics dominate and hence display Kolmogorov cascade. This is due to rotational effects introducing a significant number of vortices in the condensate, enabling their collective behaviour. This is possible frequencies above a critical frequency $\Omega>\Omega_{\mathrm{c}}$ [49], which in our case is $\Omega_{\mathrm{c}}\sim 0.3$. Similarly, the $k^{-1}$ scaling also appears at smaller length scales where single-vortex dynamics dominate, clearly indicating Vinen-like turbulence. In the presence of sustained rotation, we observe a more pronounced Kolmogorov cascade of $k^{-5/3}$ at smaller $k$ regions after at longer times. This phenomenon closely resembles the later stage evolution of rotating condensates, as previously studied in [50].
@@ IV Conclusion
We have numerically investigated the merging dynamics of Bose-Einstein condensates with random phases confined in harmonic traps, analyzing the impact of rotation on the process. Initially, we studied mergers without rotation for various merging times. All exhibited consistent scaling behaviour in the incompressible kinetic energy spectra. Density profiles at different merging times showed the formation of dark solitons due to interference patterns generated by condensate overlap. Notably, interference fringes closer to the minima of the final trap persisted longer than those near the boundaries. These dark solitons exhibited bending and subsequently decayed into vortex pairs via snake instabilities. Vortex-antivortex pairs eventually annihilated after multiple rebounds within the trap. However, while observing these features for all merging times ($t_{m}=4$, $6$, $10$, and $20$), the nature of the dark solitons, their instability-driven behaviour, and the phenomenon’s timeframe varied significantly. Increasing the merging time resulted in a decrease in the depth and size of the central dark soliton. Since soliton depth influences phase excitations, slower merging led to the attainment of the BEC’s ground state. This depth decrease was confirmed by analyzing the time-dependent variation of the forcing term. Furthermore, the angular momentum, vortex number, and incompressible kinetic energy density profiles provided temporal evidence of vortex formation and annihilation. This process occurs relatively quickly for faster mergers. The compressible kinetic energy, indicating sound waves from phase excitations, also showed a decreasing trend and almost disappeared for longer merging times. All energy profiles displayed oscillations with frequency twice that of the trap ($2\omega$) due to phase excitations. These naturally subsided and disappeared for slower mergers. In the beginning, the incompressible kinetic energy spectra exhibited $k^{-5/3}$ but only at length scales where single vortex dynamics dominate, hence ruling out the Kolmogorov cascade. Later, for a short period, they transitioned to a $k^{-1}$ scaling attributed to vortex rings formed from soliton breaking via the snake instability  [58]. However, the absence of forcing in the non-rotating case prevented these scaling laws from persisting beyond vortex annihilation. The formation of vortices through merging BECs with random initial phases, even in the absence of rotation, agrees with conclusions drawn from Refs. [26, 4].
@@
With rotation, the merging condensates display contrasting behaviour for identical merging times, consistent with observations without rotation. The evolution of the density indicates that the condensates merge off the $x$-axis despite the model restricting translational motion strictly along the $x$-axis, which results in them experiencing each other’s rotational effects as they approach. It significantly weakens the resultant dark solitons and orients with an angle to the $x$-axis. These solitons also appear to interact with the pre-existing and newly formed vortices. After merging, their rotating frequencies combine, resulting in the final condensate rotating at twice the initial frequency of the individual condensates. The significantly lower amplitude of the time-dependent forcing term confirms the weak nature of the solitons compared to the non-rotating case. The angular momentum and incompressible kinetic energy profiles indicate continuous vortex production due to sustained rotation and compressible turbulence generation. The number of vortices exceeds the anti-vortex number due to the rotation’s impact on the vortex population.
@@
The incompressible kinetic energy spectra exhibit a scaling behaviour similar to that of the non-rotating case. However, these spectra display a forcing peak at $k<2\pi/R$. Initially, the scaling range extends to length scales longer than the intervortex length $k<2\pi/\ell_{0}$ indicating the onset of Kolmogorov cascade and at later times displays Vinen-like turbulence. With sustained rotation over longer durations, we observe spectra similar to those presented in our previous work for the $\Omega=1$ case [50], where the Kolmogorov cascade is observed at smaller $k$ values.
@@
We have also carried out the simulations at various rotating frequencies and identified trends. At lower frequencies, both the scaling duration and ranges are significantly shorter, while the depth of the dark soliton is larger compared to higher rotation frequencies. Additionally, off-axis merging is less pronounced at lower frequencies. Despite these trends, the rotational case does not exhibit fundamental differences across various rotating frequencies to the same extent as the non-rotating case. Rotational merging presents interesting dynamics, particularly with large initial separations. In the case of off-axis motion, substantial separations can even lead to condensates spiralling into each other for faster merging times despite the model strictly moving the condensates along the $x$-axis.
The present study is also relevant for understanding the remnants of binary neutron-star merger event GW170817 [8, 9]. Theory predicts the aftermath of such an event should produce a stellar black hole, though observations fail to confirm this prediction conclusively. However, the following works in Refs. [10, 11], pulsars are most likely born with embryonic incompressible gluon-quark superfluid cores that become more massive as they evolve in cosmic times. During the merger, the low-energy entropy-free cores are predicted to fuse smoothly together, whereas the merger of surrounding envelopes of compressible and dissipative normal matter goes violently. Hence, the study may pave the way to explore the nature of gravitationally bound rotating BECs.
@@ Appendix A Compressible kinetic energies
Figure 16: Time evolution of compressible energy density for non-rotating condensates merging at $t_{m}=4$.
Figure 17: Time evolution of compressible energy density for rotating condensates merging at $t_{m}=4$.
Upon observing the time evolution of the forcing term, we notice that the forcing amplitude decays relatively quickly for the non-rotating case compared to the rotating one. Additionally, the forcing experienced in the condensate manifests as compressible energy generated in the merged condensate. We highlight this by analyzing the spatial density profiles of compressible kinetic energy. For the non-rotating case, we observe that the dark solitons contribute to the compressible energy density [see Fig. 16(a)]. Eventually, the energy becomes more evenly distributed and finally granulates at longer times [see Fig. 16(b)-(d)]. The decreasing amplitude of the energy density and the initial accumulation of compressible energy in solitons correlates with the evolution of the forcing term. The energy density for the rotating case further confirms this correlation, as the energy amplitude is significantly lower and distributed over a wider soliton. Similar to the forcing amplitude decreasing at a slower rate, the compressible energy peaks also do not decay quite as fast as in the non-rotating case within the same time interval (see Fig. 17). It is also noteworthy that, compared to the non-rotating case, the compressible energy is distributed as strands and is more clumped at longer times. This spatial profile resembles the long-term evolution of compressible kinetic energy shown in rotational quantum turbulence [50].

### v3

@@ III.2 Effect of rotation on the dynamics and turbulence state of the merged condensate
Figure 16: Evolution of different components of condensate energies with respect to time for (a) $t_{m}=4$, (b) $t_{m}=6$, (c) $t_{m}=10$, and (d) $t_{m}=20$ for rotating case with $\Omega_{0}=0.45$.
@@
Similar to the non-rotating case, all energy components (except for rotation energy) exhibit oscillations consistent with the trap frequency for shorter merging times. However, unlike the non-rotating case, the amplitude of these oscillations damps out quickly [see Figs. 16(b)-(d)]. The kinetic energy decreases with increasing merging time, but unlike the non-rotating merger, it does not go to zero for the $t_{m}=10$ and $t_{m}=20$ cases.
@@
Figure 17: Evolution of chemical potential (averaged over $3000$ time units from the time of merger) of the merged condensate with respect to different merging times for (a) $\Omega_{0}=0.0$ (black line with empty squares) and $\Omega_{0}=0.2$, (red line with empty circles) (b) $\Omega_{0}=0.3$ (green line with empty triangles) and $\Omega_{0}=0.45$ (blue line with empty triangles) frequencies, respectively. The dashed horizontal lines represent the ground state chemical potential for the merged condensate rotating at $2\Omega_{0}$.
@@
Figure 18: Time-averaged incompressible kinetic energy spectra for the rotating case ($\Omega_{0}=0.45$) for different merging time (a,b) for $t_{m}=4$ and (c,d) for $t_{m}=6$. (a) represents the averaged spectra in the range of time interval $t=4$ to $t=24$ and (b) $t=24$ to $t=44$. For merging time $t_{m}=6$ in the time range (c) $t=6$ to $t=26$ and (d) $t=26$ to $t=46$. Both cases initially showing $k^{-5/3}$ and later a $k^{-1}$ scaling.
@@
Achieving an adiabatic merger in the condensate depends on both the merging time ($t_{m}$) and the rotation frequency ($\Omega_{0}$), as shown in Fig. 17. Condensates rotating at frequency $\Omega_{0}$ merge to form a condensate rotating at $2\Omega_{0}$. Sufficiently long merging times ensure that this merged condensate reaches its ground state. Notably, for both non-rotating condensates and those at low rotation frequencies [black line with empty squares in Fig. 17(a)], the ground state is achieved with merging times of $t_{m}\sim 20$. In Figs. 17(b), as the rotation frequency increases, longer merging times are required for the final condensate to approach its ground state. Even without soliton decay, trap movement can create compressible excitations. These excitations persist longer at higher rotation frequencies, as rotational forcing prevents them from dissipating, thus requiring much slower mergers to suppress them.
@@
The incompressible kinetic energy spectra for the rotating case exhibit similar scaling behaviour to that of the non-rotating case (see Fig. 18). It is important to note that a forcing peak emerges due to rotation at $k<2\pi/R$ gradually diminishes for longer merging times. For cases with $t_{m}=4$ and $t_{m}=6$, a short-lived $k^{-5/3}$ scaling is initially observed, which later transforms into a $k^{-1}$ scaling, corresponding to the formation of vortex-rings (see Ref. Barenghi et al., 2023). These vortex rings are formed in the soliton decay process through snake instability, as illustrated in Ref. Verma, Rapol, and Nath, 2017.
@@
In contrast to the nonrotating case, where no scaling is observed for $t_{m}=10$ and $t_{m}=20$, the rotating case exhibits $k^{-5/3}$ scaling, albeit for a relatively shorter $k$ range initially. However, the $k^{-1}$ scaling persists for the $t_{m}=10$ case but not for the $t_{m}=20$ case, as no soliton is formed to decay into vortex-rings. In the non-rotating scenario, the scaling behaviour of the incompressible spectra disappears due to vortex annihilation. Also, in contrast to the non-rotating case, the $k^{-5/3}$ scaling appears consistently for length scales larger than inter-vortex distance $\ell_{0}$, where collective vortex dynamics dominate and hence display Kolmogorov cascade. This is due to rotational effects that introduce a significant number of vortices into the condensate, enabling their collective behaviour. This is possible for frequencies above critical frequency $\Omega_{0}>\Omega_{\mathrm{c}}$umar et al. (2019), which in our case is $\Omega_{\mathrm{c}}\sim 0.3$. Similarly, the $k^{-1}$ scaling also appears at smaller length scales where single-vortex dynamics dominate, clearly indicating Vinen-like turbulence. In the presence of sustained rotation, we observe a more pronounced Kolmogorov cascade of $k^{-5/3}$ at smaller $k$ regions after longer times. This phenomenon closely resembles the later stage evolution of rotating condensates, as previously studied in Ref. Sivakumar et al., 2024.
@@ IV Conclusion
We have numerically investigated the merging dynamics of Bose-Einstein condensates with random phases confined in harmonic traps, analyzing the impact of rotation on the process. Initially, we studied mergers without rotation for various merging times. All exhibited consistent scaling behaviour in the incompressible kinetic energy spectra. The density profiles at different merging times showed the formation of dark solitons due to interference patterns generated by condensate overlap. In particular, interference fringes closer to the minima of the final trap persisted longer than those near the boundaries. These dark solitons exhibited bending and subsequently decayed into vortex pairs via snake instabilities. Vortex-antivortex pairs eventually annihilated after multiple rebounds within the trap. However, while observing these features for all merging times ($t_{m}=4$, $6$, $10$, and $20$), the nature of the dark solitons, their instability-driven behaviour, and the phenomenon’s timeframe varied significantly. Increasing the merging time resulted in a decrease in the depth and size of the central dark soliton. Since soliton depth influences phase excitations, slower merging led to the attainment of the BEC’s ground state. This depth decrease was confirmed by analyzing the time-dependent variation of the forcing term. Furthermore, the vortex number and incompressible kinetic energy density profiles provided temporal evidence of vortex formation and annihilation. The compressible kinetic energy, indicating sound waves from phase excitations, also showed a decreasing trend and almost disappeared for longer merging times. In the beginning, the incompressible kinetic energy spectra exhibited $k^{-5/3}$ but only at length scales where single vortex dynamics dominate, hence ruling out the Kolmogorov cascade. Later, for a short period, they transitioned to a $k^{-1}$ scaling attributed to vortex rings formed from soliton breaking via the snake instabilityBarenghi et al. (2023). However, the absence of forcing in the nonrotating case prevented these scaling laws from persisting beyond vortex annihilation. The formation of vortices through merging BECs with random initial phases, even in the absence of rotation, agrees with conclusions drawn from Refs. Buchmann, Nikolopoulos, and Lambropoulos, 2009; Scherer et al., 2007. In the case of both rotating and non-rotating mergers, the interaction, potential, and kinetic energy components execute oscillations consistent with frequency $2.0\omega$. Such oscillations (generated by oscillating the traps externally) are shown to create turbulence via solitons and vortex pair annihilationsMiddleton-Spencer et al. (2023), similar to our results presented here where these oscillations are triggered by fast collisions. Conversely, for slower mergers, these oscillations get subsided, thus reducing the turbulent effects. These effects are corroborated by the scaling laws observed in the energy spectra for shorter merging times and the absence thereof for longer merging times.
@@
With rotation, the merging condensates display contrasting behaviour for identical merging times, consistent with observations without rotation. The evolution of the density indicates that the condensates merge off the $x$-axis despite the model restricting translational motion strictly along the $x$-axis, which results in them experiencing each other’s rotational effects as they approach. It significantly weakens the resultant dark solitons and orients with an angle to the $x$-axis. These solitons also appear to interact with the preexisting and newly formed vortices. After merging, their rotating frequencies combine, resulting in the final condensate rotating at twice the initial frequency of the individual condensates The significantly lower amplitude of the time-dependent forcing term confirms the weak nature of the solitons compared to the non-rotating case. The incompressible kinetic energy profiles indicate continuous vortex production due to sustained rotation and compressible turbulence generation. The number of vortices exceeds the anti-vortex number due to the rotation’s impact on the vortex population.
@@
The incompressible kinetic energy spectra exhibit a scaling behaviour similar to that of the non-rotating case. However, these spectra display a forcing peak at $k<2\pi/R$. Initially, the scaling range extends to length scales longer than the intervortex length $k<2\pi/\ell_{0}$ indicating the onset of the Kolmogorov cascade and at later times displays Vinen-like turbulence. With sustained rotation over a longer duration, we observe spectra similar to those presented in our previous work for the $\Omega_{f}=1$ caseSivakumar et al. (2024), where the Kolmogorov cascade is observed at smaller $k$ values. The presence of rotation significantly enhances vortex production, affecting turbulence in the merged condensate. For rotation frequencies below $0.3$, although the incompressible spectra show a $k^{-5/3}$ scaling, this scaling range exists only for wavenumbers larger than the intervortex distance ($2\pi/\ell_{0}$). This indicates the absence of a Kolmogorov energy cascade due to the insufficient number of vortices. Only with stronger rotation (frequencies above $0.3$) and increased vortex production do we observe scaling for wavenumbers below $2\pi/\ell_{0}$, that exhibit Kolmogorov-like cascade.
On the behaviour of the energy oscillations for the rotating case, it is quite evident that they are not as strong as those observed with the non-rotating case and decay relatively quickly for longer merging times. So, the distribution of the compressible energy is not granulated like in the non-rotating case (see. Appendix figure A.1 but appear more clumped in the form of strands (see. Appendix figure A.2). Despite weakened compressible turbulence, the incompressible kinetic energy spectra show a stronger energy cascade relative to its non-rotating counterpart, which one could attribute to the vortex production augmented due to the presence of rotation.
Barenghi et al.Barenghi et al. (2023) have highlighted the complex dynamics of quantum turbulence into three primary types: Kolmogorov turbulence, Vinen quantum turbulence, and strong quantum turbulence based on their underlying mechanisms and comparisons with classical turbulence. In our work, we observed that rapid mergers can induce turbulent instability, causing the condensate to oscillate and nucleate solitons that migrate to the periphery, reflecting and colliding to form vortex-antivortex pairs. These collisions produce sound waves and, in conjunction with density fluctuations, contribute to intricate, turbulent fluid dynamics. Although rotational forcing generates fewer sound waves, it stimulates vortex production and breakdown, resulting in a Kolmogorov cascade characteristic of strong quantum turbulenceMiddleton-Spencer et al. (2023); Barenghi et al. (2023). As vortex pairs annihilate and solitons decay, the system transitions to a Vinen turbulence regime.
@@
We have also carried out the simulations at various rotating frequencies and identified trends. At lower frequencies, both the scaling duration and ranges are significantly shorter, while the depth of the dark soliton is larger compared to higher rotation frequencies. Additionally, off-axis merging is less pronounced at lower frequencies. Despite these trends, the rotational case does not exhibit fundamental differences across various rotating frequencies to the same extent as the non-rotating case.
Finally, it is worthwhile to discuss the experimental feasibility of verifying the instability and turbulence observed in this study. Given the trap confinement parameters described in Sec. II, condensates of approximately $1\mu\mbox{m}$ size are expected to merge within milliseconds, potentially leading to turbulent behaviour. While rotational mergers are challenging to tackle due to the spiralling nature of the condensates, experimental verification of turbulence has been achieved for non-rotating superfluidsEscartín et al. (2019). One possible approach to triggering mergers is by gradually reducing the central barrier in a double-well potential, causing the condensates to combine into a single well. However, this method is not suitable for rotational mergers because the condensates collide off the merging axis. To overcome this issue, an additional velocity could be imparted to the condensates along their individual rotational motion around the $z$-axis.
@@ Appendix A Compressible kinetic energies
Figure A.1: Plots illustrating the time evolution of compressible energy density for non-rotating condensates with merging time $t_{m}=4$ at different instants of time: (a) $t=4$, (b) $t=20$, (c) $t=200$, and (d) $t=2000$.
Upon observing the time evolution of the forcing term, we notice that the forcing amplitude decays relatively quickly for the non-rotating case compared to the rotating one. Additionally, the forcing experienced in the condensate manifests as compressible energy generated in the merged condensate. We highlight this by analyzing the spatial density profiles of compressible kinetic energy. For the non-rotating case, we observe that the dark solitons contribute to the compressible energy density [see Fig. A.1(a)]. Eventually, the energy becomes more evenly distributed and finally granulates at longer times [see Fig. A.1(b)-(d)]. The decreasing amplitude of the energy density and the initial accumulation of compressible energy in solitons correlates with the evolution of the forcing term. The energy density for the rotating case further confirms this correlation, as the energy amplitude is significantly lower and distributed over a wider soliton. Similar to the forcing amplitude decreasing at a slower rate, the compressible energy peaks also do not decay quite as fast as in the non-rotating case within the same time interval (see Fig. A.2).
Figure A.2: Plots showing the time evolution of compressible energy density for rotating condensates with merging time $t_{m}=4$ at different instants of time: (a) $t=4$, (b) $t=20$, (c) $t=200$, and (d) $t=2000$.
It is also noteworthy that, compared to the non-rotating case, the compressible energy is distributed as strands and is more clumped at longer times. This spatial profile resembles the long-term evolution of compressible kinetic energy shown in rotational quantum turbulenceSivakumar et al. (2024).